# High-Coherence Kerr-cat qubit in 2D architecture

### v1

@@ I Introduction
Figure 1: Kerr-cat qubit encoding and 2D chip architecture. (a) The Bloch sphere for the Kerr-cat qubit is spanned by the cat states along the x-axis and the coherent states along the z-axis. (b) The energy pseudopotential is a double well potential at ±$\pm\alpha$ with energy gap $4K\alpha^{2}$ separating the computational space from the excited states. The Wigner function of the cat states and the excited states are shown according to the parity. (c) The circuit diagram shows the nonlinear oscillator with the ports used for control and readout. The numerical values of the design parameters are reported in Table 1. (d) The distribution of the cat states with an average photon number equal to 4 (i.e., $|\alpha|^{2}=\langle\hat{a}^{\dagger}\hat{a}\rangle$) in terms of the Fock states showing the parity of the cat states and the tail of the distribution. (e) The 2D chip (10x10 mm2) includes two uncoupled SNAILmons (yellow), each with its dedicated readout resonator and Purcell filter (light and dark green). Resonant drives at frequency $\omega_{\text{Q}}$ are delivered through weakly coupled control lines (orange), while the strong off-resonant drives used for stabilization ($\omega_{\text{S}}$) and readout ($\omega_{\text{{CQR}}}$) are delivered through a strongly coupled port with a band-block filter shown in purple. (f) The nonlinearity of the oscillator comes from two SNAIL loops biased around $\Phi_{\mathrm{ext}}/\Phi_{\mathrm{0}}=0.3$ by an off-chip flux coil. (g) Each stub of the filter consists of a meander section to have destructive interference for frequencies inside the band of the filter
@@
In particular, the family of cat qubits uses this nonlinearity to stabilize the superposition of macroscopically distinct so-called Schrödinger cat states. Depending on the details of the encoding scheme, those qubits exhibit either exponential protection or parity changes with single-photon loss, making them a valuable resource for implementing quantum error correction schemes [13, 14, 15, 16, 17, 18, 19, 20, 21]. One approach to stabilize those states with autonomous protection from bit-flips is by engineering two-photon addition and subtraction in the case of the dissipative cat. This approach offers exponential protections with increasing cat size but requires external ancillas to engineer the dissipation and the perform readout [19, 22, 23, 24].
An alternative approach to stabilize the cat states is by subjecting a Kerr nonlinear oscillator to a coherent two-photon drive [25, 26]. This approach has the advantage of creating a relatively large energy gap (5-100 MHz) between the computational space and the rest of the Hilbert space, enabling fast logic operations (20-$400$ ns) and high-fidelity readout with no added ancillas. This qubit gained a lot of attention in terms of examining its noise-biased behavior both theoretically [27, 28, 29, 30, 31] and experimentally [20, 32, 33, 34, 35]. Also, the Kerr-cat qubit makes a promising candidate to implement the XZZX surface code provided it can be realized in a scalable architecture. [36, 37, 38].
The protection against single-photon loss in this qubit can be intuitively understood by noting that the coherent states |±$\ket{\pm\alpha}$ are quasi-orthogonal eigenstates of the annihilation operator ($\hat{a}$) with eigenvalues ±$\pm\alpha$, so the dominant process of single-photon loss only leads to a linear increase in the qubit dephasing at the benefit of exponential suppression in the relaxation rate. The first realizations of this qubit were in 3D architectures and relied on strong microwave drives which led to notable heating and degraded the qubit lifetime [20, 32, 35]. Therefore, it is still an open question to see whether possible to realize large cat sizes with high coherence in 2D chips while maintaining high-fidelity operations. This uncertainty arises because the 3D environment is known to result in longer lifetimes but weaker light-matter coupling due to the larger spread of the electromagnetic field in the extra dimension [39, 40, 41].
@@
In this work, we overcome this challenge by using an effective band-block filter that enables strong coupling off-resonance (at the frequency of the stabilization drive $\omega_{\text{S}}$ and the cat readout drive $\omega_{\text{CQR}}$) while suppressing the qubit Purcell decay time orders of magnitude larger than the limit imposed by materials defects. Operating at a Kerr nonlinearity of 1.2 MHz, we observe improvements in the bit-flip time with increasing drive strength up to cat of size 10 photons. The bit-flip time exceeds 1 ms while only sacrificing the phase-flip time linearly with the cat size. Also, the large energy gap at this Kerr value enabled us to perform the highest fidelity readout and gates on Kerr-cats to date. Our work, therefore, provides a path for scaling noise-protected qubits toward fault-tolerant quantum computers.
@@ II Encoding
This Hamiltonian is now traditionally realized by charge pumping a SNAIL (Superconducting Nonlinear Asymmetric Inductive eLement) oscillator with a coherent drive at approximately twice the oscillator (i.e. qubit) frequency $\omega_{\text{S}}\approx 2\omega_{\text{Q}}$ [20, 32]. Utilizing the third-order nonlinearity of the SNAIL $\hat{\varphi}^{3}$ (where $\hat{\varphi}$ is the superconducting phase operator), a photon from the drive is consumed to create two photons in the oscillator (${a}^{\dagger 2}$) at the rate $\epsilon_{2}$; this is accompanied by the complementary process of eliminating two photons from the oscillator ($\hat{a}^{2}$) and creating a pump photon [43, 44]. The other nonlinearity of the SNAIL is a fourth-order ($\hat{\varphi}^{4}$), which in the rotating wave approximation (RWA) adds the Kerr nonlinearity $-K\hat{a}^{\dagger 2}\hat{a}^{2}$ to the Hamiltonian. We refer to this element in the Fock encoding as the SNAILmon (i.e., transmon with small anharmonicity provided by the SNAIL).
@@ III Design
Figure 3: Readout and gates with the Kerr-cat qubit. (a) The sequence to measure the QNDness of the cat quadrature readout involves two consecutive measurements to determine the probability of getting the same outcome. The demodulated output signal for the second measurement is shown in the IQ plane conditional on the first measurement being |±$\ket{\pm\alpha}$. The data is normalized with respect to the standard deviation of the Gaussian distribution ($\sigma$) of each output. For every initial state we average $10^{7}$ measurements. (b) The calibration of the phase detuning $X(\pi/2)$ gate is based on using two $X(\pi/2)$ pulses to flip the qubit and change the sign of $\innerproduct{Z}{Z}$. This gate is implemented by deforming the double-well potential, allowing the two coherent states to tunnel. The presence of the derivative in Eq.2 enables one to detune the phase in either direction. The calibration consists of finding the best choice of gate time $T_{g}$ and maximum modulation depth $\delta_{0}$ to maximize the population transfer between the coherent states. The dashed circle indicates the calibration point. (c) The calibration of the drive phase to implement the Z gate is determined by observing the fastest oscillations in $\innerproduct{Y}{Y}$ (plotted relative to the phase of $\alpha$). When driving in phase, we get an enhancement factor $2\alpha$ while driving out of phase leads to exponentially suppressed rates.
The band-block filter we use to couple strongly to the stabilization and readout drives is shown in Fig. 1(e, g). This filter consists of multiple stubs, each leading to destructive interference of the microwave signal around the qubit frequency. This is achieved by adjusting the length of the meandering section (i.e., the quarter wave resonators) and simulating the scattering parameters in Ansys-HFSS [47]. Ideally, this has high isolation at the qubit frequency ($\omega_{\text{Q}}/2\pi=5.9$ GHz) with zero attenuation at the frequency of the stabilization drive ($\omega_{\text{S}}/2\pi=11.8$ GHz) and the frequency of the cat quadrature readout ($\omega_{\text{CQR}}/2\pi=1.2$ GHz). In Fig. 2, we present an independent measurement of the transmission coefficient $\bar{S}_{21}$ of the filter in a dedicated chip (for more details about the setup see Appendix G). In Fig. 2(a), we show the behavior of normalized scattering parameter $\bar{S}_{21}$ for the full range of interest. The 30 dB of isolation at the qubit frequency corresponds to a three-order of magnitude improvement in the lifetime, pushing the Purcell limit through the strongly coupled port to 12 ms, which is much higher than the measured single photon decay time of the Fock qubit of $T_{1}=38.5$ $\mu\text{s}$. Material defects in the interfaces have been shown to limit the lifetime of 2D architectures at this range. [48, 49, 50, 51, 52, 53].
@@ IV Qubit Control and Readout
Universal single-qubit control typically requires driving on resonance to do rotations along two orthogonal axes, depending on the phase of the drive [12, 25]. However, this protocol in the cat encoding leads to an enhancement factor $2\alpha$ in the speed of the $Z(\theta)$ rotations and an exponential suppression of the $Y(\theta)$ rotations [20](For more details, check Appendix D). To overcome this challenge, it is sufficient to use the $Z(\theta)$ rotations with $X(\pi/2)$ gates to have universal control. The first proposal to do an $X(\pi/2)$ gate was demonstrated in [20] by abruptly turning off the stabilization drive for a period $\pi/2K=216$ ns. However, this protocol removes protection the stabilization drive adds to the qubit, and it tends to have worse performance for large cats [32].
@@
Fig. 3(b) shows the expectation value of Pauli Z after implementing two $X(\pi/2)$ pulses as a function of the maximum detuning $\delta_{0}$ in the phase modulation waveform and the gate time $T_{g}$. This gate implementation is resource-efficient since it does not require additional tones and does not disturb the qubit parameters compared to a fast DC flux. The coherent tunneling between the states inside the two energy wells using a controlled time-dependent detuning to perform logic gates was first examined theoretically [26, 63, 25], and the static detuning was shown to improve the lifetime experimentally [35, 32]. The simulation of this gate according to Eq. 2 is shown in Fig. 6 (more details in Appendix C.2). From the chevron plot, we chose $T_{g}=320$ ns and a detuning $\delta_{0}/K=-8.2$ for maximum population transfer.
Next, we calibrate the phase of the drive to implement $Z(\theta)$ rotations by driving on resonance. The added term to the Kerr-cat Hamiltonian has the form $\approx 2\alpha|\Omega_{\rm z}|\cos(\theta_{\rm z})\hat{Z}-2\alpha\Omega_{\rm z}\sin(\theta_{\rm z})e^{-2|\alpha|^{2}}\hat{Y}$ as long as the drive amplitude $\Omega_{\rm z}$ is small compared to the energy gap $E_{\text{gap}}/\hbar$ [27, 20]. Fig3(e) shows the protocol to calibrate the phase of the drive ($\text{Arg}(\Omega_{\rm z})=\theta_{\rm z}$) by leveraging the cat quadrature readout and the $X(\pi/2)$ gate introduced earlier. The second $X(\pi/2)$ gate enables us to measure Rabi oscillations in $\innerproduct{Y}{Y}$ as we sweep the phase of the Rabi drive $\theta_{\rm z}$ with maximum contrast $|\innerproduct{Y}{Y}|\approx 0.9$. When using the traditional $X(\pi/2)$ gate where the stabilization drive is turned off [20], the maximum contrast drops to $|\innerproduct{Y}{Y}|\approx 0.5$. We can see in Fig. 3(c) that driving in phase leads to the maximum Rabi rate while driving around $\theta_{\rm z}=\pm\pi/2$ leads to suppressed rates. For a cat of size 4, we choose a gate time of $120$ ns to implement the $Z(\pi/2)$ gate. With the $X(\pi/2)$ and the $Z(\theta)$ calibrated, we have a universal set of single-qubit gates. Using process tomography, we measure gates fidelities $\mathcal{F}_{X_{\pi/2}}=90.8\%$ and $\mathcal{F}_{Z_{\pi/2}}=91.7\%$. For the identity, we measure $\mathcal{F}_{I}=92.7\%$, which implies that state preparation errors are the limiting factor. The Pauli transfer matrices for those processes with the details of the analysis are provided in Appendix E.
@@ V Cat Qubit Lifetime
We start by measuring the lifetime of the cat qubit as a function of the cat size $\alpha^{2}=\epsilon_{2}/K$ by varying the strength of the two-photon drive $\epsilon_{2}$ as shown in Fig. 4. Every sequence begins with a readout signal to initialize the qubit in the Z-basis. To measure the lifetime of the cat states $T_{\rm C}$ (i.e., phase-flip time), we use the Kerr $X(\pi/2)$ gate introduced in [20], since it does not depend on the cat size in the calibration. For every cat size, we vary the delay time $\Delta t$ to observe the exponential decay of the coherence and perform another $X(\pi/2)$ since the cat readout projects the qubit to the Z basis. The measured lifetime in Fig. 4(a) follows the trend in grey, which is the linear trade-off for large cats described by $T_{\rm C}=T_{1}/2\langle\bar{n}\rangle$ with $\langle\bar{n}\rangle=|\alpha|^{2}(1+e^{-4|\alpha|^{2}})/(1-e^{-4|\alpha|^{2}})$ is the time-averaged number of photons in the qubit [32].
@@
The linear increase in the phase-flip rate of the cat states should come at the benefit of exponential protection against the single photon loss in the bit-flip rate. In Fig. 4(b), we examine the lifetime of the coherent states $T_{\alpha}$ (i.e. bit-flip time) as a function of the cat size. The staircase shape we observe implies that thermal excitations limit the lifetime. A bit-flip error can happen through tunneling to the higher energy states, which become asymptotically degenerate at the points where the lifetime increases sharply with the cat size. This trend was first observed in [32]. The staircase behavior is captured by heating effects in the master equation like $\kappa_{\omega_{d}/2}n_{\omega_{d}/2}\mathcal{D}\{\hat{a}^{\dagger}\}$ while the exact location of the plateaus in this model depends on the spectrum of the dephasing noise $\kappa_{\phi}(\omega)$ [64] (see Appendix B for more details). Our data up to cats of size 8 can be understood in terms of the single and multi-photon heating and dephasing of the SNAIL with thermal population $n_{\omega_{d}/2}=2\%$. This value is the lowest compared to other Kerr-cat experiments or similar strongly driven qubits[42]. However, this is still higher than the upper bound of $0.3\%$ value we measure from the residual population of the |$\ket{n=1}$ state in the Fock basis [65]. The details of the master equation model are presented in Appendix B.
@@
Next, we focus on cats of size 4, which offer the most gain in $T_{\alpha}$ without sacrificing $T_{\rm C}$ much. It is also possible to red-tune the stabilization drive to boost the lifetime of the coherent states by adding the harmonic term $\Delta\hat{a}^{\dagger}\hat{a}$ with $\Delta=\omega_{\text{Q}}-\omega_{\text{S}}/2$ to the Kerr-cat Hamiltonian in Eq.1 [32, 35]. Fig. 4(c) shows the improvement in $T_{\alpha}$ for $\epsilon_{2}/K=4$ as we sweep the detuning. The lifetime peaks around even multiples of K due to the Hamiltonian picking extra degeneracies in the excited states [35]. However, in the case of the detuned cat, the ground state of the Fock qubit |$\ket{n=0}$ does not map to the even cat state of the detuned cat qubit, so instead, we keep the stabilization drive for a time $\tau_{\text{th}}\approx T_{1}$ necessary to put the detuned cat in a thermal state, then we use the cat quadrature readout to initialize the qubit. Those results demonstrate the potential of the Kerr-cat qubit to benefit from the suppressed bit-flip rate with only a linear increase in the phase-flip rate while maintaining the ability to perform high-fidelity operations.

### v2

@@ I Introduction
Figure 1: Kerr-cat qubit encoding and 2D chip architecture. (a) The Bloch sphere for the Kerr-cat qubit is spanned by the cat states along the x-axis and the coherent states along the z-axis. (b) The energy pseudopotential is a double well potential at ±$\pm\alpha$ with energy gap $4K\alpha^{2}$ separating the computational space from the excited states. The Wigner function of the cat states and the excited states are shown according to the parity. (c) The circuit diagram shows the nonlinear oscillator with the ports used for stabilization ($c_{p}$), control ($c^{\prime}$), and readout. The numerical values of the design parameters are reported in Table 1. (d) The distribution of the cat states with an average photon number equal to 4 (i.e., $|\alpha|^{2}=\langle\hat{a}^{\dagger}\hat{a}\rangle$) in terms of the Fock states showing the parity of the cat states and the tail of the distribution. (e) The 2D chip (10x10 mm2) includes two uncoupled SNAILmons (yellow), each with its dedicated readout resonator and Purcell filter (light and dark green). Resonant drives at frequency $\omega_{\text{Q}}$ are delivered through weakly coupled control lines (orange), while the strong off-resonant drives used for stabilization ($\omega_{\text{S}}$) and cat readout ($\omega_{\text{{CQR}}}$) are delivered through a strongly coupled port with a band-block filter (purple). (f) The nonlinearity of the oscillator comes from two SNAIL loops biased around $\Phi_{\mathrm{ext}}/\Phi_{\mathrm{0}}=0.3$ by an off-chip flux coil. (g) Each stub of the filter consists of a meander section to have destructive interference around the qubit frequency.
@@
In particular, the family of cat qubits uses this nonlinearity to stabilize the superposition of macroscopically distinct so-called Schrödinger cat states. Depending on the details of the encoding scheme, those qubits can exhibit either exponential protection or parity changes with single-photon loss, making them a valuable resource for implementing quantum error correction schemes [13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24]. One approach to stabilize those states with autonomous protection from bit-flips is by engineering two-photon addition and subtraction in the case of the dissipative cat. This approach offers exponential protection with increasing cat size but requires external ancillas to engineer the dissipation and perform readout [19, 25, 26, 27].
An alternative approach to stabilize the cat states is by subjecting a Kerr nonlinear oscillator to a coherent two-photon drive [28, 29]. This approach has the advantage of creating a relatively large energy gap (5-100 MHz) between the computational space and the rest of the Hilbert space, enabling fast logic operations (20-$400$ ns) and high-fidelity readout with no added ancillas. The Kerr-cat qubit gained a lot of attention in terms of examining its noise-biased behavior both theoretically [30, 31, 32, 33, 34] and experimentally [20, 35, 36, 37, 38]. Also, this qubit makes a promising candidate to implement the XZZX surface code provided it can be realized in a scalable architecture [39, 40, 41].
The protection against single-photon loss in this qubit can be intuitively understood by noting that the coherent states |±$\ket{\pm\alpha}$ are quasi-orthogonal eigenstates of the annihilation operator ($\hat{a}$) with eigenvalues ±$\pm\alpha$, so the dominant process of single-photon loss only leads to a linear increase in the qubit dephasing at the benefit of exponential suppression in the relaxation rate. The first realizations of this qubit were in 3D architectures and relied on strong microwave drives which led to notable heating and degraded the qubit lifetime [20, 35, 38]. Therefore, it is still an open question to see whether possible to realize large cat states with high coherence in 2D chips while maintaining high-fidelity operations. This uncertainty arises because the 3D environment is known to result in longer lifetimes but weaker light-matter coupling due to the larger spread of the electromagnetic field in the extra dimension [42, 43, 44].
@@
In this work, we overcome this challenge by using an effective band-block filter that enables strong coupling off-resonance (at the frequency of the stabilization drive $\omega_{\text{S}}$ and the cat readout drive $\omega_{\text{CQR}}$) while suppressing the qubit Purcell decay rate orders of magnitude smaller than the limit imposed by materials defects. Operating at a Kerr nonlinearity of 1.2 MHz, we observe improvements in the bit-flip time with increasing drive strength up to cat of size 10 photons. The bit-flip time exceeds 1 ms while only increasing the phase-flip rate linearly with the cat size. Also, the large energy gap at this Kerr value enabled us to perform the highest fidelity readout and gates on Kerr-cats to date. Our work, therefore, provides a path for scaling noise-protected qubits toward fault-tolerant quantum computers.
@@ II Encoding
This Hamiltonian is now traditionally realized by charge pumping a SNAIL (Superconducting Nonlinear Asymmetric Inductive eLement) oscillator with a coherent drive at approximately twice the oscillator (i.e. qubit) frequency $\omega_{\text{S}}\approx 2\omega_{\text{Q}}$ [20, 35]. Utilizing the third-order nonlinearity of the SNAIL $\hat{\varphi}^{3}$ (where $\hat{\varphi}$ is the superconducting phase operator), a photon from the drive is consumed to create two photons in the oscillator (${a}^{\dagger 2}$) at the rate $\epsilon_{2}$; this is accompanied by the complementary process of eliminating two photons from the oscillator ($\hat{a}^{2}$) and creating a pump photon [46, 47, 48]. The other nonlinearity of the SNAIL is a fourth-order ($\hat{\varphi}^{4}$), which in the rotating wave approximation (RWA) adds the Kerr nonlinearity $-K\hat{a}^{\dagger 2}\hat{a}^{2}$ to the Hamiltonian. We refer to this element in the Fock encoding as the SNAILmon (i.e., transmon with small anharmonicity provided by the SNAIL).
@@ III Design
Figure 3: Readout and gates with the Kerr-cat qubit. (a) The sequence to measure the QNDness of the cat quadrature readout involves two consecutive measurements to determine the probability of getting the same outcome. The demodulated output signal for the second measurement is shown in the IQ plane conditional on the first measurement being |±$\ket{\pm\alpha}$. The data is normalized with respect to the standard deviation of the Gaussian distribution ($\sigma$) of each output. For every initial state we average $10^{7}$ measurements. (b) The calibration of the detuning $X(\pi/2)$ gate is based on using two $X(\pi/2)$ pulses to flip the qubit and change the sign of $Z\right>$. This gate is implemented by deforming the double-well potential, allowing the two coherent states to tunnel. The presence of the derivative in Eq.2 enables one to detune the phase in either direction. The calibration consists of finding the best choice of gate time $T_{g}$ and maximum modulation depth $\delta_{0}$ to maximize the population transfer between the coherent states. The dashed circle indicates the calibration point. (c) The calibration of the drive phase to implement the Z gate is determined by observing the fastest oscillations in $Y\right>$ (plotted relative to the phase of $\alpha$). When driving in phase, we get an enhancement factor $2\alpha$ while driving out of phase leads to exponentially suppressed rates.
The band-block filter we use to couple strongly to the stabilization and readout drives is shown in Fig. 1(e, g). This filter consists of multiple stubs, each leading to destructive interference of the microwave signal around the qubit frequency. This is achieved by adjusting the length of the meandering section (i.e., the quarter wave resonators) and simulating the scattering parameters in Ansys-HFSS [51]. Ideally, this filter has high isolation at the qubit frequency ($\omega_{\text{Q}}/2\pi=5.9$ GHz) with zero attenuation at the frequency of the stabilization drive ($\omega_{\text{S}}/2\pi=11.8$ GHz) and the frequency of the cat quadrature readout ($\omega_{\text{CQR}}/2\pi=1.2$ GHz). In Fig. 2, we present an independent measurement of the transmission coefficient $\bar{S}_{21}$ of the filter in a dedicated chip (for more details about the setup see Appendix G). In Fig. 2(a), we show the behavior of the normalized scattering parameter $\bar{S}_{21}$ for the full range of interest. The 30 dB of isolation at the qubit frequency corresponds to a three-order of magnitude improvement in the lifetime, pushing the Purcell limit through the strongly coupled port to 12 ms, which is much higher than the measured single photon decay time of the Fock qubit of $T_{1}=38.5$ $\mu\text{s}$. Material defects in the interfaces have been shown to limit the lifetime of 2D architectures at this range. [52, 53, 54, 55, 56, 57].
@@ IV Qubit Control and Readout
Universal single-qubit control typically requires driving on resonance to do rotations along two orthogonal axes, depending on the phase of the drive [12, 28]. However, this protocol in the cat encoding leads to an enhancement factor $2\alpha$ in the speed of the $Z(\theta)$ rotations and an exponential suppression of the $Y(\theta)$ rotations [20](For more details, check Appendix D). To overcome this challenge, it is sufficient to use the $Z(\theta)$ rotations with $X(\pi/2)$ gates to have universal control. The first proposal to do an $X(\pi/2)$ gate was demonstrated in [20] by abruptly turning off the stabilization drive for a period $\pi/2K=216$ ns. However, this protocol removes the protection the stabilization drive adds to the qubit, and it tends to have worse performance for large cats [35].
@@
Fig. 3(b) shows the expectation value of Pauli Z after implementing two $X(\pi/2)$ pulses as a function of the maximum detuning $\delta_{0}$ in the phase modulation waveform and the gate time $T_{g}$. This gate implementation is resource-efficient since it does not require additional tones and does not disturb the SNAIL parameters compared to a fast DC flux. The coherent tunneling between the states inside the two energy wells using a controlled time-dependent detuning to perform logic gates was first examined theoretically [29, 67, 28], and the static detuning was shown to improve the lifetime experimentally [38, 35]. The simulation of this gate according to Eq. 2 is shown in Fig. 6 (more details in Appendix C.2). From the chevron plot, we chose $T_{g}=320$ ns and a detuning $\delta_{0}/K=-8.2$ for maximum population transfer.
Next, we calibrate the phase of the drive to implement $Z(\theta)$ rotations by driving on resonance. The added term to the Kerr-cat Hamiltonian has the form $\approx 2\alpha|\Omega_{\rm z}|\cos(\theta_{\rm z})\hat{Z}-2\alpha\Omega_{\rm z}\sin(\theta_{\rm z})e^{-2|\alpha|^{2}}\hat{Y}$ as long as the drive amplitude $\Omega_{\rm z}$ is small compared to the energy gap $E_{\text{gap}}/\hbar$ [30, 20]. Fig. 3(e) shows the protocol to calibrate the phase of the drive ($\text{Arg}(\Omega_{\rm z})=\theta_{\rm z}$) by leveraging the cat quadrature readout and the $X(\pi/2)$ gate introduced earlier. The second $X(\pi/2)$ gate enables us to measure Rabi oscillations in $Y\right>$ as we sweep the phase of the Rabi drive $\theta_{\rm z}$ with maximum contrast |$Y\right>|\approx 0.9$. When using the traditional $X(\pi/2)$ gate where the stabilization drive is turned off [20], the maximum contrast drops to |$Y\right>|\approx 0.5$. We can see in Fig. 3(c) that driving in phase leads to the maximum Rabi rate while driving around $\theta_{\rm z}=\pm\pi/2$ leads to suppressed rates. For a cat of size 4, we choose a gate time of $120$ ns to implement the $Z(\pi/2)$ gate. With the $X(\pi/2)$ and the $Z(\theta)$ calibrated, we have a universal set of single-qubit gates. Using process tomography, we measure gates fidelities $\mathcal{F}_{X_{\pi/2}}=90.8\%$ and $\mathcal{F}_{Z_{\pi/2}}=91.7\%$. For the identity, we measure $\mathcal{F}_{I}=92.7\%$, which implies that state preparation errors are the limiting factor. The Pauli transfer matrices for those processes with the details of the analysis are provided in Appendix E.
@@ V Cat Qubit Lifetime
We start by measuring the lifetime of the cat qubit as a function of the cat size $\alpha^{2}=\epsilon_{2}/K$ by varying the strength of the two-photon drive $\epsilon_{2}$ as shown in Fig. 4. Every sequence begins with a readout signal to initialize the qubit in the Z-basis. To measure the lifetime of the cat states $T_{\rm C}$ (i.e., phase-flip time), we use the Kerr $X(\pi/2)$ gate introduced in [20], since it does not depend on the cat size in the calibration. For every cat size, we vary the delay time $\Delta t$ to observe the exponential decay of the coherence and perform another $X(\pi/2)$ since the cat readout projects the qubit to the Z-basis. The measured lifetime in Fig. 4(a) follows the trend in grey, which is the linear trade-off for large cats described by $T_{\rm C}=T_{1}/2\langle\bar{n}\rangle$ with $\langle\bar{n}\rangle=|\alpha|^{2}(1+e^{-4|\alpha|^{2}})/(1-e^{-4|\alpha|^{2}})$ the time-averaged number of photons in the qubit [35].
@@
The linear increase in the phase-flip rate of the cat states should come at the benefit of exponential protection against the single photon loss in the bit-flip rate. In Fig. 4(b), we examine the lifetime of the coherent states $T_{\alpha}$ (i.e. bit-flip time) as a function of the cat size. The staircase shape we observe implies that thermal excitations limit the lifetime. A bit-flip error can happen through tunneling to the higher energy states, which starts to become asymptotically degenerate at the points where the lifetime increases sharply with the cat size. This trend was first observed in [35]. The staircase behavior is captured by heating effects in the master equation like $\kappa_{\omega_{d}/2}n_{\omega_{d}/2}\mathcal{D}\{\hat{a}^{\dagger}\}$ while the exact location of the plateaus in this model depends on the spectrum of the dephasing noise $\kappa_{\phi}(\omega)$ [68] (see Appendix B for more details). Our data up to cats of size 8 can be understood in terms of the single and multi-photon heating and dephasing of the SNAIL with thermal population $n_{\omega_{d}/2}=2\%$. This value is the lowest compared to other Kerr-cat experiments or similar strongly driven qubits[45]. However, this is still higher than the upper bound of $0.3\%$ value we measure from the residual population of the |$\ket{n=1}$ state in the Fock basis [69]. The details of the master equation model are presented in Appendix B.
@@
Next, we focus on cats of size 4, which offer the most gain in $T_{\alpha}$ without sacrificing $T_{\rm C}$ much. It is also possible to red-tune the stabilization drive to boost the lifetime of the coherent states by adding the harmonic term $\Delta\hat{a}^{\dagger}\hat{a}$ with $\Delta=\omega_{\text{Q}}-\omega_{\text{S}}/2$ to the Kerr-cat Hamiltonian in Eq.1 [35, 38]. Fig. 4(c) shows the improvement in $T_{\alpha}$ for $\epsilon_{2}/K=4$ as we sweep the detuning. The lifetime peaks around even multiples of $K$ due to the Hamiltonian picking extra degeneracies in the excited states [38]. However, in the case of the detuned cat, the ground state of the Fock qubit |$\ket{n=0}$ does not map to the even cat state of the detuned cat qubit, so instead, we keep the stabilization drive for a time $\tau_{\text{th}}\approx T_{1}$ necessary to put the detuned cat in a thermal state, then we use the cat quadrature readout to initialize the qubit. Those results demonstrate the potential of the Kerr-cat qubit to benefit from the suppressed bit-flip rate with only a linear increase in the phase-flip rate with increasing cat size while maintaining high-fidelity operations.